# When Unsupervised Domain Adaptation meets One-class Anomaly Detection: Addressing the Two-fold Unsupervised Curse by Leveraging Anomaly Scarcity

### v2

@@ 6.1 Experimental Setting
We evaluate our approach on four standard UDA benchmark datasets, Office-Home [35], Office31 [35],VisDA [28], and PACS [20]. The types of domain shift of each dataset is described in Section A of the Supplementary. For the AD task, we adopt a standard one-vs-all protocol, where a single class is considered normal and the remaining are anomalies. We adopt the experimental protocol of previous DA works [22, 43, 7] to allow for a fair comparison –that is, we show results on ten classes from the ClipArt and Product domains for Office-Home, ten classes from Webcam and Amazon for Office31, and twelve classes from the domains of Computer Aided Designs (CAD) (synthetic objects) and real object photos of VisDA. On PACS, like [4], we consider the Photo domain as source and the remaining three domains as targets.
@@
As no other works on UDA for semantic image UAD were previously introduced, we compare our method with several few-shot SoA approaches. More specifically, we consider BiOST [7] which is a one-shot approach and TSA [21], ILDR [19], IRAD [43], and MsRA [22] that are few-shot methods. Furthermore, we introduce a few-shot variant (Ours-Few-shot) of our approach, which augments the target domain with normal and pseudo-anomalous samples similar to [11]. This augmentation yields semantically positive and negative pairs [11], which are useful for the contrastive alignment strategy described in Section 5. Further details are given in Section A of the Supplementary.
@@
In all experiments, the source set has only normal data, while the unlabeled target set includes mostly normals with 10% randomly sampled anomalies. Training uses SGD with a cosine-annealing scheduler, learning rate of $10^{-3}$, weight decay of $5\times 10^{-7}$, batch size 256 and $\lambda_{1}$ and $\lambda_{2}$ are set to 1. CLIP-ViT-B32 is the frozen visual encoder $\psi$ for feature clustering. To align with the setting of the baselines [22, 43, 7, 21], ResNet50 is the trainable backbone $f$, initialized on ImageNet [9]. K-means clustering [14] uses 2, 10, and 5 components for Office, VisDA, and PACS, respectively. Like the baselines, the few-shot experiments use 10 (Office, PACS) and 100 shots (VisDA) labeled as normal, respectively.
@@
Our method outperforms previous SoA on all benchmarks of our evaluation, as shown in Table 1 and Table 2. More specifically, our fully unsupervised UDA-UAD importantly improves upon previous few-shot SoA on C $\rightarrow$ P and W $\rightarrow$ A of the Office-Home [41] and Office31 [35] datasets. In addition, we observe an improvement of over $10\%$ in the VisDA dataset [28] with the fully unsupervised methodology over previous few-shot approaches despite being challenged by the two-fold unsupervised curse. These results highlight the relevance of the proposed method, even in the presence of a large domain gap, as in the case of synthetic CAD images and real-world photos.
@@
Furthermore, we compare the performance of our model to two pretrained visual encoders [30], namely ResNet50 and CLIP-ViT-B32 in Table 3. While the CLIP-ViT-B32 architecture achieves an average AUC of 72.08%, our unsupervised method (75.54%) still outperforms it on the VisDA dataset [28]. In contrast, the ResNet50 shows a significantly lower performance, with an average AUC of only 53.45%. These results demonstrate that despite their strong performance, pretrained visual encoders are not specifically tailored for the domain adaptation task; thus, they remain vulnerable to domain shift. Therefore, training domain adaptation-specific models is still necessary to effectively bridge the gap between two given domains.
@@ 6.3 Additional Experiments
Unless stated otherwise, all the following experiments are performed on VisDA [28]. Additional results on different components of our method are given in the Supplementary.
@@
The results presented in Table 3 compare pretrained visual encoders and source-only detectors with different adaptation paradigms, i.e., few-shot, unsupervised, and supervised (oracle). The source-only finetuned model improves slightly over the pretrained ResNet50 visual encoder [15] but still has lower performance than the adaptation approaches, achieving an average AUC of 62.84%. Among the few-shot methods, our few-shot variant outperforms BiOST [7] and MsRA [22], achieving the highest AUC of 71.93%, which is comparable to the performance of a pretrained CLIP-ViT-B32 visual encoder. However, our unsupervised adaptation method surpasses all these models, with an average AUC of 75.54% indicating its ability to effectively mitigate domain gaps without relying on labeled target data. This can be explained by the fact that after clustering, our model has access to more representative normal target data than few-shot models, hence better generalizing to the target normal class. On the other hand, the Oracle, which has access to the target labels, achieves the highest performance (81.65%). The small gap between our unsupervised method and the oracle demonstrates the effectiveness of our approach even without supervision.
@@
$\pm$±std $\pm$19.57 $\pm$17.83 $\pm$8.55 $\pm$7.55 $\pm$5.79 $\pm$9.08 $\pm$9.80 $\pm$6.11
Table 3: AUC (%) performance on the target domain of our UDA anomaly detector on VisDA [28] compared with various adaptation paradigms (from zero-shot, i.e., pretrained Visual encoders, few-shot, to supervised, i.e., Oracle).
@@
Table 4 provides the results obtained when each component, namely the adaptation loss, the dominant cluster identification through clustering, the use of an auxiliary visual encoder $\psi(\mathcal{D}^{t})$ or the trainable features $f(\mathcal{D}^{t})$. The results show that without adaptation, a model trained only on source data generalizes poorly to the target domain with only 62.84%. Direct adaptation of the source and the unlabeled target without clustering leads to inconsistent results, indicating low generalization capabilities to the target domain. Introducing clustering results in a significant performance boost. This can be seen when clustering is applied to the original representations of the feature extractor, as the performance improves by +5.63%, highlighting the importance of identifying the dominant cluster prior to alignment. Note that our method still outperforms the best few-shot baseline MsRA [22] (68.47% vs 67.28%) with just clustering and alignment (i.e. w/o CLIP) across all the VisDA classes. Finally, the best results are achieved when all components are combined. This setup boosts the average AUC to 75.54% on all VisDA classes. The substantial performance gains can be attributed to CLIP’s rich visual features, which, together with clustering and alignment, help achieve a more robust anomaly detector capable of better handling domain shift.
@@
VisDA 62.84$\pm$08.55 75.54$\pm$10.23 72.24$\pm$08.81 74.14$\pm$07.44 71.65$\pm$06.68
[…]
Table 5: Ablation on different clustering techniques. GMM and K-means use $10$ components for VisDA and $2$ for other datasets. k=2 for VisDA and k=1 for the remaining datasets.
Clustering methods.
Figure 5: K-Means[14] components variation on VisDA.
$f$: ResNet50 + $\psi$: CLIP-ViT-B32
DSVDD [33] MSC [31]
ZS Src UDA ZS Src UDA
53.45 62.84 75.54 74.39 72.87 81.10
(+12.7$\uparrow$) (+8.23$\uparrow$)
Table 6: Our UDA approach on two anomaly detection methods: DSVDD [33] and Mean-shifted Contrastive Learning (MSC) [31]. ZS and Src mean Zero-shot and Source only, respectively.
Beyond DSVDD.
To assess whether our alignment approach applies to other unsupervised AD methods, we replace DSVDD with Mean-shifted Contrastive loss (MSC) [31] in Table 6. MSC adapts contrastive loss to the one-class setting by shifting augmented representations of the normal samples toward the mean of pretrained normal features, preserving their compactness. Similar to [32, 6], it uses a kNN density estimator to detect anomalies. Our results suggest that both methods benefit from the adaptation, as a consistent average improvement of +12.7% and +8.23% is seen across all the 12 classes of VisDA.
Alignment strategies.
By comparing several alignment strategies in Table 7, we observe that any alignment strategy, in general, improves the performance consistently for all adaptation benchmarks. Contrastive alignment consistently outperforms other adaptation losses, including statistical (MMD) [12] and adversarial (GRL) [10] strategies.
Dataset Source Only GRL [10] MMD [12] Contrastive [26]
VisDA 62.84$\pm$08.55 71.84$\pm$10.89 73.12$\pm$10.36 75.54$\pm$9.80
@@
Table 8 compares the results of GNL [4] with our proposed UDA method on the PACS dataset [20], with Photo as the source and Art, Cartoon, and Sketch as the target domains. It can be seen that our UDA approach consistently outperforms GNL [4], particularly on Cartoon and Sketch domains. This suggests that unlike DG methods, which aim to generalize to any unseen domain solely by training on the source domain, UDA can be more effective for semantic UAD since it exposes the model to the target domain during training, even if it is unlabeled.
@@
DG GNL [4] 65.62 67.96 62.39
DA MsRA [22] (Few-shot) 71.43 69.89 61.87
@@ 7 Limitations and Future Work
Our method, presented in Section 5, is one possible solution for addressing the problem of UDA for UAD. However, it is worth noting that it was tested in the context of semantic anomaly detection [38], adopting a one-vs-all protocol, to facilitate the comparison with the closest baselines, namely [22, 43]. These methods typically require the use of global features in contrast to standard anomaly detection, where fine-grained representations are usually targeted. For that reason, our method focuses mostly on global representations, while local features would be conceptually more suitable for fine-grained anomaly detection. In future works, we aim to extend our study to fine-grained anomaly detection by exploiting more relevant local representations.

### v3

@@ 6.1 Experimental Setting
We evaluate our approach on four standard UDA benchmark datasets, Office-Home [office31], Office31 [office31],VisDA [visda-2017], and PACS [pacs-2017]. The types of domain shift of each dataset is described in Section A of the Supplementary. For the AD task, we adopt a standard one-vs-all protocol, where a single class is considered normal and the remaining are anomalies. We adopt the experimental protocol of previous DA works [msra-2021, irad-2023, biost-2019] to allow for a fair comparison –that is, we show results on ten classes from the ClipArt and Product domains for Office-Home, ten classes from Webcam and Amazon for Office31, and twelve classes from the domains of Computer Aided Designs (CAD) (synthetic objects) and real object photos of VisDA. On PACS, like [ad-shift-2023], we consider the Photo domain as source and the remaining three domains as targets.
@@
As no other works on UDA for semantic image UAD were previously introduced, we compare our method with several few-shot SoA approaches. More specifically, we consider BiOST [biost-2019] which is a one-shot approach and TSA [tsa-2021], ILDR [ildr-2019], IRAD [irad-2023], and MsRA [msra-2021] that are few-shot methods. Furthermore, we introduce a few-shot variant (Ours-Few-shot) of our approach, which augments the target domain with normal and pseudo-anomalous samples similar to [jigsaw-neg-pairs-2021]. This augmentation yields semantically positive and negative pairs [jigsaw-neg-pairs-2021], which are useful for the contrastive alignment strategy described in Section 5. Further details are given in Section A of the Supplementary.
@@
In all experiments, the source set has only normal data, while the unlabeled target set includes mostly normals with 10% randomly sampled anomalies. Training uses SGD with a cosine-annealing scheduler, learning rate of $10^{-3}$, weight decay of $5\times 10^{-7}$, batch size 256 and $\lambda_{1}$ and $\lambda_{2}$ are set to 1. CLIP-ViT-B32 is the frozen visual encoder $\psi$ for feature clustering. To align with the setting of the baselines [msra-2021, irad-2023, biost-2019, tsa-2021], ResNet50 is the trainable backbone $f$, initialized on ImageNet [imagenet-2009]. K-means clustering [kmeans-1979] uses 2, 10, and 5 components for Office, VisDA, and PACS, respectively. Like the baselines, the few-shot experiments use 10 (Office, PACS) and 100 shots (VisDA) labeled as normal, respectively.
@@
Our method outperforms previous SoA on all benchmarks of our evaluation, as shown in Table 1 and Table 2. More specifically, our fully unsupervised UDA-UAD importantly improves upon previous few-shot SoA on C $\rightarrow$ P and W $\rightarrow$ A of the Office-Home [officehome] and Office31 [office31] datasets. In addition, we observe an improvement of over $10\%$ in the VisDA dataset [visda-2017] with the fully unsupervised methodology over previous few-shot approaches despite being challenged by the two-fold unsupervised curse. These results highlight the relevance of the proposed method, even in the presence of a large domain gap, as in the case of synthetic CAD images and real-world photos.
@@
Furthermore, we compare the performance of our model to two pretrained visual encoders [clip-2021], namely ResNet50 and CLIP-ViT-B32 in Table 3. While the CLIP-ViT-B32 architecture achieves an average AUC of 72.08%, our unsupervised method (75.54%) still outperforms it on the VisDA dataset [visda-2017]. In contrast, the ResNet50 shows a significantly lower performance, with an average AUC of only 53.45%. These results demonstrate that despite their strong performance, pretrained visual encoders are not specifically tailored for the domain adaptation task; thus, they remain vulnerable to domain shift. Therefore, training domain adaptation-specific models is still necessary to effectively bridge the gap between two given domains.
@@ 6.3 Additional Experiments
Unless stated otherwise, all the following experiments are performed on VisDA [visda-2017]. Additional results on different components of our method are given in the Supplementary.
@@
The results presented in Table 3 compare pretrained visual encoders and source-only detectors with different adaptation paradigms, i.e., few-shot, unsupervised, and supervised (oracle). The source-only finetuned model improves slightly over the pretrained ResNet50 visual encoder [resnet-2016] but still has lower performance than the adaptation approaches, achieving an average AUC of 62.84%. Among the few-shot methods, our few-shot variant outperforms BiOST [biost-2019] and MsRA [msra-2021], achieving the highest AUC of 71.93%, which is comparable to the performance of a pretrained CLIP-ViT-B32 visual encoder. However, our unsupervised adaptation method surpasses all these models, with an average AUC of 75.54% indicating its ability to effectively mitigate domain gaps without relying on labeled target data. This can be explained by the fact that after clustering, our model has access to more representative normal target data than few-shot models, hence better generalizing to the target normal class. On the other hand, the Oracle, which has access to the target labels, achieves the highest performance (81.65%). The small gap between our unsupervised method and the oracle demonstrates the effectiveness of our approach even without supervision.
@@
[…]
Table 3: AUC (%) performance on the target domain of our UDA anomaly detector on VisDA [visda-2017] compared with various adaptation paradigms (from zero-shot, i.e., pretrained Visual encoders, few-shot, to supervised, i.e., Oracle).
@@
Table 4 provides the results obtained when each component, namely the adaptation loss, the dominant cluster identification through clustering, the use of an auxiliary visual encoder $\psi(\mathcal{D}^{t})$ or the trainable features $f(\mathcal{D}^{t})$. The results show that without adaptation, a model trained only on source data generalizes poorly to the target domain with only 62.84%. Direct adaptation of the source and the unlabeled target without clustering leads to inconsistent results, indicating low generalization capabilities to the target domain. Introducing clustering results in a significant performance boost. This can be seen when clustering is applied to the original representations of the feature extractor, as the performance improves by +5.63%, highlighting the importance of identifying the dominant cluster prior to alignment. Note that our method still outperforms the best few-shot baseline MsRA [msra-2021] (68.47% vs 67.28%) with just clustering and alignment (i.e. w/o CLIP) across all the VisDA classes. Finally, the best results are achieved when all components are combined. This setup boosts the average AUC to 75.54% on all VisDA classes. The substantial performance gains can be attributed to CLIP’s rich visual features, which, together with clustering and alignment, help achieve a more robust anomaly detector capable of better handling domain shift.
@@
Table 5: Ablation on different clustering techniques. GMM and K-means use 1010 components for VisDA and 22 for other datasets. k=2 for VisDA and k=1 for the remaining datasets.
Clustering methods.
Figure 5: K-Means[kmeans-1979] components variation on VisDA.
$f$: ResNet50 + $\psi$: CLIP-ViT-B32
DSVDD [dsvdd-2018] MSC [cmeanshift-2023]
ZS Src UDA ZS Src UDA
53.45 62.84 75.54 74.39 72.87 81.10
[…]
Table 6: Our UDA approach on two anomaly detection methods: DSVDD [dsvdd-2018] and Mean-shifted Contrastive Learning (MSC) [cmeanshift-2023]. ZS and Src mean Zero-shot and Source only, respectively.
Beyond DSVDD.
To assess whether our alignment approach applies to other unsupervised AD methods, we replace DSVDD with Mean-shifted Contrastive loss (MSC) [cmeanshift-2023] in Table 6. MSC adapts contrastive loss to the one-class setting by shifting augmented representations of the normal samples toward the mean of pretrained normal features, preserving their compactness. Similar to [panda-2021, red-pandas], it uses a kNN density estimator to detect anomalies. Our results suggest that both methods benefit from the adaptation, as a consistent average improvement of +12.7% and +8.23% is seen across all the 12 classes of VisDA.
Alignment strategies.
By comparing several alignment strategies in Table 7, we observe that any alignment strategy, in general, improves the performance consistently for all adaptation benchmarks. Contrastive alignment consistently outperforms other adaptation losses, including statistical (MMD) [mmd-2006] and adversarial (GRL) [grl-2015] strategies.
Dataset Source Only GRL [grl-2015] MMD [mmd-2006] Contrastive [infonce-loss-2018]
[…]
Table 7: Performance in terms of AUC ($\%$) using different adaptation losses on the three UDA benchmarks.
@@
Table 8 compares the results of GNL [ad-shift-2023] with our proposed UDA method on the PACS dataset [pacs-2017], with Photo as the source and Art, Cartoon, and Sketch as the target domains. It can be seen that our UDA approach consistently outperforms GNL [ad-shift-2023], particularly on Cartoon and Sketch domains. This suggests that unlike DG methods, which aim to generalize to any unseen domain solely by training on the source domain, UDA can be more effective for semantic UAD since it exposes the model to the target domain during training, even if it is unlabeled.
@@
DG GNL [ad-shift-2023] 65.62 67.96 62.39
DA MsRA [msra-2021] (Few-shot) 71.43 69.89 61.87
@@ 7 Limitations and Future Work
Our method, presented in Section 5, is one possible solution for addressing the problem of UDA for UAD. However, it is worth noting that it was tested in the context of semantic anomaly detection [semantic-ad-2024], adopting a one-vs-all protocol, to facilitate the comparison with the closest baselines, namely [msra-2021, irad-2023]. These methods typically require the use of global features in contrast to standard anomaly detection, where fine-grained representations are usually targeted. For that reason, our method focuses mostly on global representations, while local features would be conceptually more suitable for fine-grained anomaly detection. In future works, we aim to extend our study to fine-grained anomaly detection by exploiting more relevant local representations.
@@ 8 Conclusion
9 Acknowledgments
This work was supported by the Fonds National de la Recherche (FNR) under the grant agreements no. BRIDGES2021/IS/16353350/FaKeDeTeR and UNFAKE, ref. 16763798, and by Post Luxembourg.